If I am doing a RICE table and the concentrations of the compounds are given at equilibrium, will it ALWAYS follow the coefficients from the chemical equation?

This is a conceptual question. It does not have an actual table.
Answer #1

Yes, it have to always follow the coeficients from BALANCED CHEMICAL EQUATION.

Reason.

Chemical reaction always follows STOICHIOMETRY.

we will illustrate this by taking an general example

let us take

Ag) = 20 (9) + 3D (9)

Where,

A is reactant

C and D are products

1,2,3 are stichiometric coefficients of A,C and D respectively

let us assume the above equation is balanced

Equation indicate that one mole of A on reaction gives 2 moles of C and 3 moles of D.

Remember in reaction moles may or may not be conserved but mass is always conserved (means total mass of reatants is equal to the total mass products)

Hence while making ICE table we have to t consideration of coefficients

ICE is a simple acronym for the titles of the first column of the table.

  • I stands for initial concentration.
  • C stands for the change in concentration.
  • E is for the concentration when the reaction is at equilibrium.

      Ag) = 20 (9) + 3D (9)

Initial concentration: C 0 0

Concentration at

equilibrium : C(1- \alpha ) 2C\alpha 3C\alpha

Where ,

C is molar concentration

\alpha = degree of dissociation = (number of moles dissociated)/(Total number of moles taken)

We can also define \alpha as fraction of molecules dissociated.

we have to consider the coefficients